1. 7.23 × 1024 Al atoms = grams Al
1
Expert's answer
2022-02-07T23:45:02-0500

Solution:

"moles= \frac{atoms}{Avagadro \;number}"


Moles of Al "=\frac{7.23 \times10^{23}}{6.023 \times 10^{23}}=12 (\frac{mol}{litres})"


Weight of Al = "moles \times molecular\; weight = 12 \times 27 = 324 grams"


Answer: 324 grams of Aluminium.